# Ionic association and Wien effect in 2D confined electrolytes

### v1

@@ II.4.1 Formation of ionic clusters
In previous sections, the absence of short distance repulsion between ions weakened the ionic structure, making the distinction between free ions and pairs quite blurry. Without repulsive potential, the distance between ions in a pair fluctuates around 0. As there is no permanent dipoles in this case, their interactions with other ions are very weak. In all atom MD, where ions also interact through repulsive LJ interactions, pairs fluctuates around a finite distance (see Figure 1.(b) and Figure 4.(a)). In this case, interaction with pairs are no longer negligible, allowing the formation of more complex structure.
Figure 4: Structure in explicit solvent simulations and its relation with correlation functions (a) Screenshot from EMD simulations for $T^{*}=0.1$. We observe 2 types of direct bonding, the contact bonds (1) when there is no water between ions, and solvent separated bonds (3) when both ions are bond, with solvent in between. We also observe indirect bonding (2) between ions, which is a consequence of the presence of clusters. (b) and (d) Running coordination function as a function of the distance between the ions, respectively for a reduced Coulomb temperature of 0.4 (monovalent ions) and 0.1 (divalent ions). The red and blue curves corresponds respectively to correlation between opposite and identical ions. We also identifies some of the well defined jump, which corresponds to spike in the correlation function. The plain pink vertical line corresponds to the contact bonds. The plain pink and light blue line corresponds to multiples of the distance between contact bond, and corresponds to indirect bond in linear chains of ions with alternating sign. The dotted pink and blue vertical line corresponds to solvent separated bonds. (c) Cluster size distribution $N_{l}$ defined as the number of clusters of connected ions (ions at distances smaller than $7\text{\AA}$ from at least one other ion in the cluster). In blue and red respectively the data for a reduced Coulomb temperature of $0.4$ (monovalent ions) and $0.1$ (divalent ions).
Figure 4 shows respectively the oppositely and identically charged ions running coordination numbers $N_{\text{op}}$ and $N_{\text{id}}$, for monovalent ions ($T^{*}\sim 0.4$, panel (b)) and divalent ions ($T^{*}\sim 0.1$, panel (d)). Both display well defined jumps, corresponding to thin peaks in the correlation function i.e. stiff minima in the free energy. This is a direct consequence of the presence of repulsive interaction.
From comparative analysis between the simulations movies 4.(a) and coordination curves 4.(b) and (d), we can relate the structures and the peaks. We observe both direct bonds between opposite ions (non mediated by other ions) and indirect bonds (mediated by other ions, could not exist in an ionic pair). Most of the direct bonds are contact bonds at small distances $d\sim 2\ \text{\AA}$. We also observe some solvent separated bonds, mediated by water molecules, at larger distances $d\sim 6\ \text{\AA}$.
The position of the peaks associated to indirect bonds gives us information on the structure of the clusters. For example in our system, we mostly observe peaks at multiples of the size of contact bond, which is a signature of the predominance of linear ionic chains, in agreement with was can be seen in simulations movies.
@@
Other worksRobin, Kavokine, and Bocquet (2021); Zhao et al. (2021) that used different simulations techniques reported various possible structures of ionic clusters, depending for example on the chemical nature of ionsRobin, Kavokine, and Bocquet (2021); Zhao et al. (2021) or of the wallFong et al. (2024). Overall, the relations between short-range interaction and short-range structure seem non universal and is in practice difficult to analyse. Instead, we now focus on long-range correlations, in comparison with BD simulations.
II.4.2 Pairing transition in all-atom simulation
As in Brownian dynamics we could use transport simulation in order to study pairing. Indeed, from the definition of the free carrier density (Equation (23)):
$n_{f}^{\text{cluster}}(E)=\frac{j(E)}{j_{0}(E)},$ (45)
if we neglect any variation of the mobility of the field. However, as in BD simulations, this definition is not adapted to very small field. Instead, we can use the following geometric criterion.
We can define a cluster as an ensemble of ions that are at distance smaller than a distances $d_{c}$ to at least another ion in the cluster. $d_{c}$ is a small distance, below which we consider 2 opposite ions to be bound. We will use in the following $d_{c}=7\text{\AA}$ for every charge. This allows to take into account both the contact bonds (vertical pink lines in 4) and the solvent separated bonds (vertical dashed pink lines in 4).
The distribution $N_{l}$ of clusters of size $l$ is shown in Figure 4.(c) for a reduced Coulomb temperature of 0.4 (in blue), corresponding to monovalent ions, and 0.1 (in red), corresponding to divalent ions. We observe a global exponential decay of the cluster numbers with their size, and clusters with even number of ions are more abundant.
This can be explained by the fact that clusters with an odd number of ions are charged, as they cannot have the same number of anions and cations. However, we only observe neutral even clusters. The disparity is a consequence of the free energy cost to have a charged cluster. As expected, the disparity between odd and even clusters increase when the ionic charges is increased.
In order to quantify this effect, the free charge carrier density is extracted from MD simulations. The result is shown in Figure 3.(b). Below $T^{*}\simeq 0.25$, the free charge carrier obtained from MD simulations is close to zero, and increase above, following the same tendency of BD simulations. We can use Equation (40) to fit the simulations points, and estimate the critical temperature. The critical temperature estimated this way (green cross in Figure 3) is coherent with the theoretical prediction (green line in Figure 3). As in BD simulations, the general curve does not exactly match the theoretical curve, as a consequence of the different treatment of the short-range interaction. In addition of that, the cluster density is MD simulations might evolve with the reduced Coulomb temperature.
At small reduced Coulomb temperature, the small deviation from 0 at small temperature can be explained by the fact that the charge density stabilises on very long timescales, because of the slow dynamic of the clusters.
We can write more formally the analogy between MD and BD system from the definition of $q(r)$. We can separate the sum between ions in same (intra) or different (inter) clusters. We obtain:
$q(r)=(1-n_{f}^{\text{cluster}})q_{\text{intra}}(r)+n_{f}^{\text{cluster}}q_{\text{inter}}(r).$ (46)
Where $q_{\text{intra}}$ and $q_{\text{inter}}$ are normalised to 1. For each ion $i$ inside a cluster $c$, we can write $\vec{r}_{i}=\vec{r}_{c}+\delta\vec{r}_{i}$, where we expect $\delta\vec{r}_{i}$ to be negligible compared to the distance between clusters. Neglecting multi-polar terms yields to:
$q_{\text{inter}}(r)\sim-\left\langle\sum_{c}\sum_{c^{\prime}\neq c}\frac{z_{c}z_{c^{\prime}}}{n_{f}^{\text{cluster}}}\delta(r-|\vec{r}_{c}-\vec{r}_{c^{\prime}}|)\right\rangle.$ (47)
This corresponds to long range interactions between a density $n_{f}^{\text{cluster}}$ of objects of charge $z_{c}=\pm 1$, which is the system described by the BD simulations. In the same manner, when we increase the Coulombic interaction, the cost of creating an ionic atmosphere around a charged cluster becomes more important, dragging the system to local neutrality through the formation of neutral clusters. This yields to a very similar transition.
@@ III.1 Non-linear behavior of the ionic current
$\delta j(E)\simeq j(E)-n_{f}(0)j_{0}=\delta n_{f}(E)j_{0}.$ (48)
In terms of conductivity, this yields to the following contributions:
@@
Figure 2.b and .d show the free ion fraction as a function of the electric field below and above the transition. When $E\to 0$, we observe that the free ion fraction tends linearly to a constant above the transition $T^{*}>T_{c}^{*}$, and vanished like a power law below the transition. In the limit of large applied fields, we instead observe that $n_{f}\to 1$: we recover Ohm-like conduction, as in fully dissociated electrolytes. The corresponding ionic current $I$ and conductivity $\sigma$ are plotted in Figure 5a and b.
@@
Figure 5: Comparison between BD transport simulations and theory. (a-b) Ionic current (a) and conductivity (b) for reduced Coulomb temperature of 0.1 (red) and 0.4 (blue), compared with Ohm’s law (dashed black line). Symbols: BD simulations. Solid lines: Theoretical model, Equation (84). (c) Symbols: Effective association constant $K(E)$ computed from the free ion fraction at various electric field, for a reduced Coulomb temperature $T^{*}$ of 0.1 (red) and 0.4 (blue). Solid lines: power-law fit, allowing to determine the value of the exponent $\beta$. (d) Symbols: Exponent $\alpha$ defined by the non-linearity in the IV curve, $j(E)\propto|E|^{\alpha}$, obtained from $\beta$ through $\alpha=1-\frac{\beta}{2}$. Black solid line: Theoretical prediction of the 2D Wien effect, see Equation (86). Dotted line: bulk exponent $\alpha=2$, obtained through Onsager’s theory of the Wien effect Onsager (1934).
III.2 Determination of the exponent at low field
In order to study the response of the system, it is easier to define an effective association constant $K(E)$ through:
$K(E)=\frac{n_{p}(E)}{\delta n_{f}(E)^{2}}=\frac{1-n_{f}(E)}{\left(n_{f}(E)-n_{f}(0)\right)^{2}}.$ (50)
In Figure 5.c, we plot the effective association constant as a function of the electric field. We observe that the curves are well approximated by a power law of the external field, with an exponent $\beta$ that we obtain by fitting the curve of $K(E)$ with $E$ in logarithmic scale (Fig. 5c). At small field, we have:
$\delta j(E)\underset{E\to 0}{\propto}\frac{E}{{\sqrt{K(|E|)}}},$ (51)
hence the current also follows a power law with the external field. The absolute value underlines the fact that the effective association constant - and then the conductivity - is symmetric under field inversion. We can thus define the exponent $\alpha(T^{*})$ of the power law:
$\delta j(E)\underset{E\to 0}{\propto}{E\times|E|^{\alpha(T^{*})-1}},$ (52)
which is, as expected, antisymmetric under field inversion. This exponent can be extracted from the slope from Figure 5.c, as $\alpha=1-\beta/2$. We find that this exponent strongly depends on temperature: we plot the evolution of $\alpha$ with the reduced Coulomb temperature in Figure 5.d. We observe that the value of the exponent depends on conditions (such as the strength of the interactions), which strongly contrasts with the original second Wien effect, which predicts an universal exponent of 2 for the transport of weak electrolyte regardless of details of interaction. We seem to recover this limit when the interaction between the ion becomes weak $T^{*}\to 1$ (blue curve in Figure 5.b). Surprisingly, the exponent drastically changes at a value close to the critical temperature, echoing the transition defined at equilibrium.
We checked that this exponent only depends on the strength of the interaction and not other interaction parameters, such as the short-distance cut-off $C$. The exponent obtained from our BD simulations is also in agreement with the exponents obtained in previous NEMD simulations of the system Robin, Kavokine, and Bocquet (2021).
III.3 Theory
@@
In this section, we derive analytically the expression of the non-universal exponent $\alpha(T^{*})$. Our approach is based on an extension of Onsager’s study of the second Wien effect for bulk 3D weak electrolytes Onsager (1934). The generalization to 2D has been first developed in Ref. Robin, Kavokine, and Bocquet (2021) and we elaborate on this descritpion here to compare with our numerical results. In what follows, we restrict ourselves to the case of the ideal 2D Coulomb gas ($\xi=\infty$), and focus on the paired regime, ($T^{*}<T^{*}_{c}$).
III.3.1 An Onsager’s approach of the 2D Wien effect
@@
We assume that the fraction $n_{f}$ follows a generic evolution equation, introducing a dissociation (resp. association) constant $\tau_{d}$ (resp. $\tau_{a}$):
@@
At thermal equilibrium (in the absence of external field), the ratio of these to constants is given by $e^{2\mu_{e}}$, as discussed in the previous section. When driven out of equilibrium however, these two quantities may depend on the applied electric field $\boldsymbol{\mathrm{E}}=E\,\boldsymbol{\mathrm{\hat{x}}}$ acting along $x$.
@@
The first term of the potential corresponds to the unscreened interaction of two ions in confinement. This assumption is valid if the system is sufficiently paired up ($T^{*}<T^{*}_{c}$) so that the influence of other free ions can be neglected. Quantitatively, the Debye length $\lambda_{\text{D}}$ diverges if the system is fully paired, and therefore is not a relevant length scale. The second term corresponds to the external field, characterized by the following length scale:
@@
where $C$ is a positive constant independent of $r$. In other words, $g_{a}$ describes a background of free ions recombining with the central ion to form new pairs, and $g_{d}$ pairs that break under the electric field. These two functions will allow us to compute $\tau_{a}$ and $\tau_{d}$.
@@
III.3.2 Self-similarity of the correlation function
@@
This is not the correct solution of the problem however, as the Boltzmann distribution only makes sense at thermal equilibrium. In particular, it does not verify the correct boundary conditions at $r\to\infty$. However, for $u\ll 1$ the external field is negligible compared to the field created by the central ion and pairs are in quasi-equilibrium (ions are strongly correlated and remain bounded over long timescales). Therefore, we admit that $g_{d}$ is the unique solution of the following problem:
@@
Assuming each free ion contributes linearly to conduction, we obtain the ionic current due to the Wien effect:
$j(E)\propto{E\times|E|^{1/2T^{*}}}.$ (79)
@@
In the next section, in order to explain this this behavior, we now take into account the formation of larger ionic clusters in presence of an external field.
III.3.3 Ionic clusters and anisotropic correlations
@@
$j(E)\propto E\times n_{x}(E)\propto{E\times|E|^{\alpha(T^{*})-1}},$ (85)
@@
III.3.4 Comparison with simulations and discussion
The comparison between our BD simulations and Equation (85) is plotted in Figure 5.a and b. Below the Kosterlitz-Thouless transition, the agreement is quantitative and the observed exponent $\alpha(T^{*})$ in simulations matches the theoretical prediction:
$\boxed{\text{For }T^{*}<T_{c},\quad\sigma(E)\propto|E|^{1/4T^{*}},\quad\alpha\simeq 1+\frac{1}{4T^{*}}.}$ (87)
@@ III.3.4 Comparison with simulations and discussion
$\boxed{\text{For }T^{*}>T_{c},\quad\sigma(E)-\sigma(0)\propto|E|,\quad\alpha\simeq 2,}$ (88)
@@
The above derivation was performed assuming that all ions are paired up. In what follows, we account for ions that remain free even in absence of any field by adding a contribution $n_{f}(0)$ to the above result; this term is determined by the procedure described in Section II.2.1. The above sections were dedicated to the theoretical analysis of the 2D Wien effect at low temperature ($T^{*}<T^{*}_{c}$). We find that the conductivity of 2D confined electrolytes evolves as a power law of the electric field, with a non-universal exponent $1/4T^{*}$. This contrasts with the 3D bulk case, where the Wien exponent is 1.

### v2

@@ II.2.2 Molecular dynamics simulations and effect of the short-range interactions
In previous sections, we discussed BD simulations where ions do not have any repulsion potential, meaning that short-distance structures could be adequately resolved. Without repulsive potential, the distance between ions in a pair fluctuates around 0. As there is no permanent dipole in this case, their interactions with other ions are very weak. In all atom MD, where ions also interact through repulsive LJ interactions, pairs fluctuates around a finite distance (see Figure 1.(b) and Figure 2.(b)). In this case, interaction with pairs are no longer negligible, allowing the formation of more complex structures.
@@
Other worksRobin, Kavokine, and Bocquet (2021); Zhao et al. (2021) that used different simulations techniques reported various possible structures of ionic clusters, depending for example on the chemical nature of ionsRobin, Kavokine, and Bocquet (2021); Zhao et al. (2021) or of the wallFong et al. (2024). In our case, ion correlation functions reveal a short-distance structuration due to contact and solvent-separated ion binding, up to clusters of size $\sim 6$ in absence of electric fields (Figure S3). Overall, the relations between short-range interaction and short-range structure seem non universal and is in practice difficult to analyse. Instead, we can focus on the free ion fraction, in comparison with BD simulations.
We find that, like in BD, the free ion fraction $n_{f}$ drops to zero below a critical temperature, as shown in Figure 2.(d) (green symbols). The behavior of $n_{f}$ with temperature is well-described by theoretical predictions; however the value of the critical temperature is found to differ from BD simulations, with $T^{*}_{c}\simeq 0.24$. This discrepancy can be attributed to two facts. First, in MD simulations, we only account for a single layer of wall atoms: the electrolyte is thus effectively confined by a slab of empty space, with $\epsilon_{r}=1$, leading to a much higher value of the dielectric length $\xi=28\,$nm. Besides, the minimal inter-ion approach distance is now unequivocally set by the physical size of ions, set by Lennard-Jones short-distance interactions. For our choice of simulation parameters, the distance of physical contact is $r_{0}=0.18\,$nm (corresponding to approximately twice the ionic radius of sodium). For these values of $\xi$ and $r_{0}$, we obtain from our theoretical model $T^{*}_{c}=0.24$, very close to the value infered from simulations (Figure 2.(d), green line).
To obtain the theoretical curves of Figure 2.(d), one must also specify the quantity denoted $\zeta$ in Eq. (12). It can be interpreted as a “volume fraction”: if one fixes the position of the first ion (assumed to be a disk of radius $r_{0}/2$), one must place the second ion exactly at a distance $r_{0}$ from the first ion’s center to form a contact pair. Within that disk of radius $r_{0}$, the pair occupies a fraction $\zeta_{\text{MD}}\simeq 0.36$ (corresponding to the central ion and around half of the second one). If instead ions are allowed to overlap, as in BD, then that fraction drops to approximatively $\zeta_{\text{BD}}\simeq 0.25$ as paired ions almost fully overlap. In practice, we treat $\zeta$ as a fitting parameter for the sake of simplicity, and obtain $\zeta_{\text{MD}}=0.33$ and $\zeta_{\text{BD}}=0.22$, very close to estimated values.
Similarly, we show in SI that the fraction of charged clusters also exhibit a drop below $T^{*}_{c}$. However, we find that the equilibration time of clusters grows rapidly as temperature is lowered: due to our limited simulation time, we find that $n_{f}^{\text{cluster}}$ is not exactly zero at very low temperature.
It should be noted that in our theory, the chemical equilibrium (Eq. (12)) should in principle be updated to take the formation of clusters into account, instead of only ion pairs. In practice, however, large cluster only represent a small fraction of ions (see Figure S3) and Eq. (12) can be used as an approximation.
@@ III.1 Non-linear behavior of the ionic current
$\delta j(E)\simeq j(E)-n_{f}(0)j_{0}=\delta n_{f}(E)j_{0},$ (26)
with $j_{0}=\sigma_{0}E$ the current density of a fully dissociated electrolyte. In terms of conductivity, this yields to the following contributions:
@@
Figure 3.a shows the conductivity as a function of the electric field below and above the transition. When $E\to 0$, we observe that the conductivity tends linearly to a constant above the transition $T^{*}>T_{c}^{*}$, and vanishes like a power law below the transition. In this regime, we will define the exponent $\alpha(T^{*})$:
$|\delta j(E)|\underset{E\to 0}{\propto}{|E|^{\alpha(T^{*})}}.$ (28)
@@
In the limit of large applied fields, we instead observe that $n_{f}\to 1$: we recover Ohm-like conduction, as in fully dissociated electrolytes. The corresponding ionic current $I$ is plotted in Figure 4.b.
@@
Figure 3: Comparison between BD transport simulations and theory. (a-b) Ionic current (a) and conductivity (b) for reduced Coulomb temperature of 0.1 (red) and 0.4 (blue), compared with Ohm’s law (dashed black line). Symbols: BD simulations. Solid lines: Theoretical model, Equation (61).
III.2 Theory
@@
In this section, we derive analytically the expression of the non-universal exponent $\alpha(T^{*})$. Our approach is based on an extension of Onsager’s study of the second Wien effect for bulk 3D weak electrolytes Onsager (1934). The generalization to 2D has been first developed in Ref. Robin, Kavokine, and Bocquet (2021) and we elaborate on this description here to compare with our numerical results. In what follows, we restrict ourselves to the case of the ideal 2D Coulomb gas ($\xi=\infty$), and focus on the paired regime, ($T^{*}<T^{*}_{c}$).
III.2.1 An Onsager’s approach of the 2D Wien effect
@@
We assume that the fraction $n_{f}$ follows a generic evolution equation, introducing a dissociation (resp. association) time $\tau_{d}$ (resp. $\tau_{a}$):
@@
At thermal equilibrium (in the absence of external field), the ratio of these two times is given by $e^{2\mu_{e}}$, as discussed in the previous section. When driven out of equilibrium however, these two quantities may depend on the applied electric field $\boldsymbol{\mathrm{E}}=E\,\boldsymbol{\mathrm{\hat{x}}}$ acting along $x$.
@@
Here, we made the approximation that $\log r/(r+\xi)\sim\log r$, which is reasonable for $r<\xi$. Since $\xi=14\,$nm, this approximation is valid over the typical lengthscales of ionic correlations. The first term of the potential corresponds to the unscreened interaction of two ions in confinement. This assumption is valid if the system is sufficiently paired up ($T^{*}<T^{*}_{c}$) so that the influence of other free ions can be neglected. Quantitatively, the Debye length $\lambda_{\text{D}}$ diverges if the system is fully paired, and therefore is not a relevant length scale. The second term corresponds to the external field, characterized by the following length scale:
@@
where $C$ is a positive constant independent of $r$ and $\rho$ is the average ionic density far from the central ion. In other words, $g_{a}$ describes a background of free ions recombining with the central ion to form new pairs, and $g_{d}$ pairs that break under the electric field. These two functions will allow us to compute $\tau_{a}$ and $\tau_{d}$.
@@
III.2.2 Self-similarity of the correlation function
@@
This is not the correct solution of the problem however, as the Boltzmann distribution only makes sense at thermal equilibrium. In particular, it does not verify the correct boundary conditions at $r\to\infty$. However, for $u\ll 1$ the external field is negligible compared to the field created by the central ion and pairs are in quasi-equilibrium (ions are strongly correlated and remain bounded over long timescales). Therefore, the total correlation function $g$ should diverge like $g_{0}$ for $u\to 0$. Since $g_{a}$ is bounded around $u=0$, this condition also provides us a boundary condition for $g_{d}$, which is therefore the solution of:
@@
It should be noted that this result is general for any system with Onsager’s pairing kinetics, and does not depend on details of ionic interactions, as long as $\tau_{a}$ and $\tau_{d}$ are known. In the limit of weak applied field, $\tau_{d}\gg\tau_{a}$ and $n_{f}\ll 1$. We obtain:
$n_{f}\simeq\sqrt{\frac{\tau_{a}}{\tau_{d}}},$ (55)
with $\tau_{a}$ and $\tau_{d}$ given by Eqs. (42) and (53) respectively. Assuming each free ion contributes linearly to conduction, $j(E)\propto E\times n_{f}(E)$ and we obtain the ionic current due to the Wien effect:
$|j(E)|\propto{|E|^{1+1/2T^{*}}}.$ (56)
@@
In the next section, in order to explain this this behavior, we now take into account the formation of larger ionic clusters in presence of an external field. We do so by first rederiving our result through scaling arguments, and then amending it using the same scaling logic.
Lastly, the above derivation critically relies on the (quasi-)logarithmic shape of the interaction potential: because its typical lengthscale $\xi$ is much larger than any other physical lengthscale in the problem, ionic correlations become self-similar. This situation is unique to the 2D case; for example, in bulk electrolytes, ions interact through the usual $1/r$ Coulomb’s law (which has a typical lengthscale set by the Bjerrum length $\ell_{\text{B}}$, see Section II.A.2) and computing correlations is much more mathematically involved Onsager (1934).
III.2.3 Scaling laws and anisotropic conduction
@@
$|j(E)|\propto|E|\times n_{x}(E)\propto{\times|E|^{\alpha(T^{*})}},$ (62)
@@
III.2.4 Comparison with simulations and discussion
Figure 4: Symbols: Exponent $\alpha$ defined by the non-linearity in the IV curve, $|j(E)|\propto|E|^{\alpha}$. Black solid line: Theoretical prediction of the 2D Wien effect, see Equation (63). Dotted line: bulk exponent $\alpha=2$, obtained through Onsager’s theory of the Wien effect Onsager (1934).
The comparison between our BD simulations and Equation (62) is plotted in Figure 4.a and b. Below the Kosterlitz-Thouless transition, the agreement is quantitative and the observed exponent $\alpha(T^{*})$ in simulations matches the theoretical prediction. Conductivity scales like:
${\text{For }T^{*}<T_{c},\quad\sigma(E)\propto|E|^{\alpha-1},\quad\alpha\simeq 1+\frac{1}{4T^{*}}.}$ (64)
@@ III.2.4 Comparison with simulations and discussion
${\text{For }T^{*}>T_{c},\quad\sigma(E)-\sigma(0)\propto|E|^{\alpha-1},\quad\alpha\simeq 2,}$ (65)
@@
The above derivation was performed assuming that all ions are paired up. In what follows, we account for ions that remain free even in absence of any field by adding a contribution $n_{f}(0)$ to the above result; this term is determined by the procedure described in SI. The above sections were dedicated to the theoretical analysis of the 2D Wien effect at low temperature ($T^{*}<T^{*}_{c}$). We find that the conductivity of 2D confined electrolytes evolves as a power law of the electric field, with a non-universal exponent $1/4T^{*}$. This contrasts with the 3D bulk case, where the Wien exponent is 1.